OpenStudy (anonymous):

The regular price of a child's entry ticket to a water park is $6 less than that for an adult's. The park offers half off all entry tickets during the off-peak season. The Jordans paid a total of $69 for 2 adult tickets and 1 child's ticket to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. :

OpenStudy (anonymous):

69 = x + one-half(x - 6) What is the regular price of a child's ticket? $38 $42 $48 $66

OpenStudy (anonymous):

help plz!

OpenStudy (anonymous):

do you think you can help?

OpenStudy (anonymous):

@FlufflePuffle

OpenStudy (flufflepuffle):

I can try.

OpenStudy (anonymous):

The regular price of a child's entry ticket to a water park is $6 less than that for an adult's. The park offers half off all entry tickets during the off-peak season. The Jordan's paid a total of $69 for 2 adult tickets and 1 child's ticket to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. :

OpenStudy (flufflepuffle):

It's 48.

OpenStudy (flufflepuffle):

Plug it in to the equation.

OpenStudy (anonymous):

k thx for your help

OpenStudy (flufflepuffle):

69=48+1/2(48-6) 69=48+1/2(42)

OpenStudy (flufflepuffle):

69=48+21

OpenStudy (flufflepuffle):

69=69

OpenStudy (anonymous):

oh now i get it ! ha

OpenStudy (anonymous):

thx agian

OpenStudy (flufflepuffle):

No problem.